Find the surface area generated by rotating the given curve about the y-axis. x = 312, y = 2, osts 4
Answer #1

CamScanner Seanned with on 1 4 + 1564 17 ) 1 4 + 1734 17 - 170817) si 148 = tlfzx98 - tlf 582 x 9 + b ] 181 - s! 9 | า tij sxS Q Find the surface ova generated by rotating the given curre about the younis . 1 x = 3t2, y = 2t3, o²tly. som 6t, dy - 672Temecues [Green - zeléd.) 2 + Juppies [ turto - & - 215+i) { 48) = fo=4+1) +- (1– 2st ) * ] U81 2 ti Lauen -tron f 1-431- C18